on Saturday Roberta swam for 5/6 in the morning and for 1/2 hours in the afternoon. how many hours did Roberta swim that day?


1
Expert's answer
2021-04-11T13:44:44-0400

on Saturday Roberta swam for 5/6 in the morning and for 1/2 hours in the afternoon. how many hours did Roberta swim that day?

We have that Roberta swam for 5/6 in the morning and for 1/2 hours in the afternoon.

Thus that day Roberta swam:


"\frac{5}{6}+\frac{1}{2}=\frac{5}{6}+\frac{3}{6}=\frac{8}{6}=\frac{4}{3}=1\frac{1}{3}"



Answer: Roberta swam 1 hour 20 minutes that day.